OpenStudy (sterling234):

Help please Solve the system of equation by elimination and substitution -66x-5y=-24 11x+4y=-21

OpenStudy (jclark):

What do you already know to do?

OpenStudy (sterling234):

I did the problem i just ended up with a fraction so i don't know if its right

OpenStudy (jclark):

fractions can do that. Show me a few of your steps

OpenStudy (triciaal):

the objective of this method is to make the coefficient of one variable equal and opposite so when the equations are added that variable is eliminated. you are left with a linear equation. solve for that variable then use the value in either of the expressions to solve for the eliminated variable. It is a very good practice to check your solution in either of the originals

OpenStudy (sterling234):

This is what i did -66x-5y=-24 11x+4y=-21 i multiplied the second equation by 6 to cancel out x so it would look like -5y=-24 24y=-126

OpenStudy (jclark):

yep

OpenStudy (sterling234):

then i got -19=-150/-19

OpenStudy (sterling234):

y*

OpenStudy (triciaal):

when you solve a system of equations it means you are finding the (x,y) that will satisfy both equations at the same time

OpenStudy (jclark):

24v + -5y = 19y so y = -150/19

OpenStudy (jclark):

it's ugly, but the math is correct

OpenStudy (sterling234):

i know i have to find the value of (x,y) that will work in both equations

OpenStudy (sterling234):

I was just working on graphing this and i didnt get the same answer thats my main problem

OpenStudy (jclark):

What have you tried to eliminate y?

OpenStudy (triciaal):

when you solve a system of equations it means you are finding the (x,y) that will satisfy both equations at the same time

OpenStudy (sterling234):

i tell you the solution i got

OpenStudy (jclark):

I would not recommend trying to find a solution by graphing...! Your first approach was fine, you did the calculations right. It's just an awkward answer

OpenStudy (jclark):

go ahead

OpenStudy (sterling234):

y=-150/19 x= 201/209

OpenStudy (sterling234):

and its a project so i sorta have to do it :/

OpenStudy (jclark):

oof. Use really big graph paper...

OpenStudy (sterling234):

But for substitution can you help me?

OpenStudy (jclark):

your solution is correct

OpenStudy (sterling234):

i have to show it in the substitution method

OpenStudy (jclark):

where do you think you should start with substitution?

OpenStudy (sterling234):

isolate one of the variables

OpenStudy (jclark):

okay, pick one

OpenStudy (sterling234):

x

OpenStudy (jclark):

first or second equation?

OpenStudy (sterling234):

what do you recommend :/

OpenStudy (jclark):

If nothing jumps out at you, which ever looks less complicated

OpenStudy (sterling234):

2nd

OpenStudy (jclark):

okay - so for now forget about the first equation, and simply solve for x

OpenStudy (sterling234):

y must be 0 right?

OpenStudy (jclark):

no, just get x by itself

OpenStudy (sterling234):

i could add 5 y to -24

OpenStudy (sterling234):

i mean -21

OpenStudy (jclark):

forget about the first equation... just focus on the second 11x+4y = -21 solve for x

OpenStudy (jclark):

is that what you were doing?

OpenStudy (sterling234):

yeah but i substituted y for 0

OpenStudy (jclark):

youre thinking about intercepts

OpenStudy (jclark):

get x by itself

OpenStudy (jclark):

you got this...

OpenStudy (sterling234):

ok 11x+4y=-21 -11 -11 x+4y=-10

OpenStudy (sterling234):

....

OpenStudy (jclark):

right idea, wrong numbers...

OpenStudy (sterling234):

so 4?

OpenStudy (jclark):

you are indeed subtracting soemthing from both sides, but not 11x

OpenStudy (jclark):

just 4?

OpenStudy (sterling234):

4y

OpenStudy (jclark):

bingo

OpenStudy (jclark):

so now you have 11x = -21 - 4y

OpenStudy (sterling234):

so 11x+4y=-21 -4y -4y 11x= -21- 4y

OpenStudy (sterling234):

Now i remember lol

OpenStudy (jclark):

correct. what's next?

OpenStudy (jclark):

:)

OpenStudy (sterling234):

substitute that into the other equation

OpenStudy (jclark):

almost. one more step to solve for x first

OpenStudy (jclark):

it isn't quite by itself yet...

OpenStudy (sterling234):

how would i get rid of y?

OpenStudy (jclark):

you have 11x = -21-4y. You want "x = blah blah blah" What needs to be done so there is no "11"?

OpenStudy (sterling234):

oh so y can remain so i can subtract 11 from -21 making it x=-31+4y

OpenStudy (jclark):

closer still... you can only subtract 11 if it's being added. It is not. It is being multiplied to x. So to undo that you must_______ by 11

OpenStudy (sterling234):

add

OpenStudy (sterling234):

sorry

OpenStudy (jclark):

no worries

OpenStudy (sterling234):

so now substitute?

OpenStudy (sterling234):

-66(-10+4y)-5y=-24 ?

OpenStudy (jclark):

you are so close. You have the substitution correct

OpenStudy (jclark):

You're just missing one thing on your solution for x Here's a simpler example: 2x + y = 5 2x = 5 - y x = (5-y) /2

OpenStudy (sterling234):

...

OpenStudy (sterling234):

So what exactly did i do wrong

OpenStudy (jclark):

11x = -21-4y divide both sides by 11

OpenStudy (jclark):

x = (-12-4y)/11

OpenStudy (jclark):

er... -21

OpenStudy (sterling234):

So adding could'nt be effective?

OpenStudy (jclark):

no, because we aren't adding or subtracting 11 and x. we are multiplying 11 by x. So the only way to eliminate the 11 is to divide it

OpenStudy (sterling234):

Ok I'll revise the equation

OpenStudy (jclark):

yea, your substitution was spot on except for the dividing by 11 part

OpenStudy (sterling234):

-66(-12-4y/11)-5y=-24

OpenStudy (jclark):

correct. and yuck. what a mess.

OpenStudy (sterling234):

I know right

OpenStudy (sterling234):

Now distribute -66 right

OpenStudy (jclark):

yes...

OpenStudy (jclark):

let me see what you have after that, I want to give you some advice, but I don't want to interupt your train of thought. you're doing great

OpenStudy (sterling234):

792+264y/11-5y=-24

OpenStudy (sterling234):

do i distribute 11 too?

OpenStudy (jclark):

no, because it isn't outside any parenthesis

OpenStudy (sterling234):

Oh ok

OpenStudy (sterling234):

just checking if i was correct

OpenStudy (jclark):

however, this is wehre the advice comes in. It would be much easier to deal with this if it werent for that fraction

OpenStudy (sterling234):

so would we multiply now to get rid of it

OpenStudy (jclark):

yes!

OpenStudy (jclark):

becareful about how you do it tho

OpenStudy (jclark):

\[\frac{ 792+ 264y}{ 11 } -5y=-24\]

OpenStudy (jclark):

you have to multiply every term by 11

OpenStudy (sterling234):

i thpought it was only on each side

OpenStudy (sterling234):

thought*